OpenStudy (anonymous):

Find the exact value of 2sin135sin75? Would appreciate any help!

OpenStudy (anonymous):

Ok, the answer is apparetnyl \[(1 + \sqrt{3})/2\], but I keep getting sqrt(6)/2?? Please help me.

OpenStudy (anonymous):

Here is my work: \[2(\sqrt{2}/2)\sin75°\] \[2(\sqrt{2}/2)(\sqrt{3}/2)\] = \[\sqrt{6}/2\] But the answer is supposed to be 1 + sqrt(3) / 2.. What am I doing wrong?

OpenStudy (mimi_x3):

Ok lemme take a look But ya better Fan+Medal+Testimonial deal?

OpenStudy (anonymous):

Sure! Thank you @Mimi_x3

OpenStudy (anonymous):

I also got \[2(\sqrt{2}/2)(\sqrt{1}/2) \] but that doesn't get me the right answer.

OpenStudy (mimi_x3):

Hey I am going to eat dinner but in the mean time maybe @dan815 Can help you. Otherwise ill be back in 20 minutes

OpenStudy (anonymous):

Thank you! :)

OpenStudy (dan815):

sin(75deg) is not root3/2

OpenStudy (anonymous):

@dan815 - I'm aware, but I just tried doing all I could and trying all possibilities. I converted sin75 into sin150/2 which got me root3/2 :(

OpenStudy (anonymous):

it went from root(1-cos150/2) = root3/2 but It's wrong. I just don't know what to do since product - to - sum includes sin + sin but this is sin*sin so I'm lost!

OpenStudy (dan815):

its not sin*75

OpenStudy (dan815):

its sin of 75

OpenStudy (dan815):

sin is a big function that 75 is put into

OpenStudy (dan815):

Sin(75) + sin(75) is not equal to sin(150)

OpenStudy (dan815):

do you half angle formula or sin addition formula

OpenStudy (dan815):

sin(a+b)=? or sin(2a)=?

OpenStudy (anonymous):

I tried sin(theta/2) and that's how I got 150 since I didn't know sin of 75.

OpenStudy (anonymous):

But I think I got the answer!! :) I did 2*1/2(cos(135-75) - cos(135+75). Which equals 1(cos60 - 210) = 1(1/2 + root3/2) = (1 + root3)/2!! I got it! :D

OpenStudy (anonymous):

Thanks for the help guys. I guess I was just stressing out and couldn't think straight. :)

OpenStudy (dan815):

ok good =] glad i cud help